OpenStudy (anonymous):

Help? Which is less than 0.65 ? A 13 -- 20 ------------ B 14 -- 25 ------------ C 4 -- 5 ------------ D 7 -- 10 (Thanks<3)

OpenStudy (anonymous):

This is easy. Just use a calculator to divide each one. :) 13 / 20 = 0.65 14 / 25 = 0.56 4 / 5 = 0.8 7 / 10 = 0.7 So the only answer less than 0.65 is B.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Oh, I see....Thank you also!!

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

But @Calcmathlete she needs to find a number LESS than 0.65, so she couldn't do it that way.

OpenStudy (anonymous):

Oh ok. I didn't see that!

OpenStudy (anonymous):

Yeah. Nevermind then!

OpenStudy (anonymous):

It's okay. :)

OpenStudy (anonymous):

But that would have been a good way to do it if it had asked for a number equal to 0.65. :)